OpenStudy (anonymous):

- 1/2 x + 2 = x + 7 I still dont understand how to do it

OpenStudy (anonymous):

how to do what?

OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

- 1/2 x + 2 = x + 7

OpenStudy (anonymous):

x==-10/3

OpenStudy (anonymous):

How did you get to that answer?

OpenStudy (anonymous):

I want to learn it, not just get answers haha

OpenStudy (radar):

Follow the "golden rule" of algebra. Thou shalt do to one side of the equal sign as thou do to the other side.

OpenStudy (radar):

In this case subtract x. You now have -1.5x + 2 = 7 Do you see why?

OpenStudy (anonymous):

Yeah

OpenStudy (radar):

Continuing, now subtract 2. You now have: -1.5x= 5 Again did you follow?

OpenStudy (jiteshmeghwal9):

\[\Huge{\frac{1}{2}x+2=x+7}\]\[\Huge{\frac{x+2}{2}=\frac{x+7}{1}}\]By cross multiplication we have,\[\Huge{(x+2)=2(x+7)}\]Solve this for \(x\).

OpenStudy (anonymous):

Yes

OpenStudy (radar):

I think you are on track. We only now have to do is divide by -1.5 We now have x = -(5/1.5) Using calculator divide 5 by 1.5 and prefix the minus sign.

OpenStudy (jiteshmeghwal9):